25) write the equation of each hyperbola.
Center(-2,-4); a=2; area of fundamental rectangle is 36 square units.
so stuck :/

Answer by lwsshak3(11628)

25) write the equation of each hyperbola.
Center(-2,-4); a=2; area of fundamental rectangle is 36 square units.
**
Assume hyperbola has a horizontal transverse axis.(can't tell if horizontal or vertical from given info)
Standard form: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center.
..
area of fundamental rectangle = 36 =2a*2b
2b=36/2a=36/4=9
b=9/2
b^2=81/4
..
Equation of given hyperbola:
(x-h)^2/a^2-(y-k)^2/b^2=1
(x+2)^2/4-4(y+4)^2/81=1